OpenStudy (anonymous):

Given the linear equation y = -3x – 2, find the y-coordinates of the points (0, ), (-3, ), (-1, ). Please show all of your work. Plot these points and graph the linear equation. Submit your graph through the dropbox.

OpenStudy (bahrom7893):

what are y values when x values are the ones that are given?

OpenStudy (anonymous):

There aren't any values given.

OpenStudy (bahrom7893):

y = -3*0-2 y = -3*(-3)-2 y = -3*(-1)-2

OpenStudy (bahrom7893):

the x values are all given

OpenStudy (anonymous):

ye4s

OpenStudy (bahrom7893):

so here are the y values: y = -2 y = 7 y=1

OpenStudy (anonymous):

So what are my coordinates?

OpenStudy (bahrom7893):

(0,-2), (-3,7), (-1,1)

OpenStudy (anonymous):

Can you show me where do they go on the graph??